OpenStudy (anonymous):

write a general formula to describe the variation: x varies directly with the cube of y; x = 8 when y = -8

OpenStudy (anonymous):

x = ky^3 substitute x = 8 and y = -8 8 = k(-8)^3 8 = -512k k = -8/512 So x = (-8/512)y^3.

OpenStudy (anonymous):

Thanks could you help with another one. Write a general formula to describe the variation: x varies inversely with the square of y; x = -1, when y = 5